What is the value of x in the proportion 32:(x+2) = 20:(25+x)?
Answer by Alan3354(69443)

What is the value of x in the proportion 32:(x+2) = 20:(25+x)?
---------------
32:(x+2) = 20:(25+x)
You can multiply means and extremes, the 2 inside terms and the 2 outside terms, and set them equal. Or, view these as fractions, then cross multiply. The result is the same:
32*(25+x) = 20*(x+2)
32x + 800 = 20x + 40
12x = -760
x = -190/3